OpenStudy (anonymous):

****************Probability***************** A bag contains five red chips, 3 blue chips, and 2 yellow chips. Two chips are selected from the bag, without replacement. Determine the probability that two chips of the same color are selected

OpenStudy (anonymous):

gotta do it for each color

jabez177 (jabez177):

Oh can I give direct answers Oh right no sorry

OpenStudy (anonymous):

i did for Red, I got 1/2 Blue, 3/10 Yellow, 1/5

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

but I don't know what to do after that

OpenStudy (anonymous):

and alright thank you

OpenStudy (anonymous):

start with yellow. That's the easiest

OpenStudy (anonymous):

two red is not right probability first is red is \(\frac{1}{2}\) then the probability second is red given the first is red is \[\frac{4}{9}\] since there are 9 left and 4 red ones so probababiltyi both are red is \[\frac{1}{2}\times \frac{4}{9}=\frac{2}{9}\]

OpenStudy (anonymous):

other two are similar once you get the other two, add them up

OpenStudy (anonymous):

oh right I forgot to take into account that the chip wasn't put back

OpenStudy (anonymous):

getting 2 blue is very similar, except you start with 3, not 5

OpenStudy (anonymous):

for blue i got 1/15, and yellow 1/45

OpenStudy (anonymous):

i didn't do it but i can check

OpenStudy (anonymous):

1/15 is right

OpenStudy (anonymous):

so is 1/45

OpenStudy (anonymous):

you said to add them up at the end so i make all the denominators the same right?

OpenStudy (anonymous):

yes of course or use a calculator but you can use a denominator of 45 for each

OpenStudy (anonymous):

14/45?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

oh alright perfect thank you you helped alot!

OpenStudy (anonymous):

yw glad to help